OpenStudy (anonymous):

The perimeter of a rectangle is 48 inches. The length is three times the width. Find the dimensions of the rectangle thanks.

OpenStudy (anonymous):

width is 6 length is 18

OpenStudy (anonymous):

3w+3w+w+w=48 8w=48 w=6

OpenStudy (anonymous):

The equation for the perimeter is 2l+2w=48 As the length is three times the width: l = 3w. We can't solve for two variables, but we have an equation for l which we can subtitute in: \(2(3w)+2w=48\) \(6w+2w=48\) \(8w=48\) \(w=6\) Then substitute w in l=3w to find length: \(l=3(6)=18\) Thus width is 6 inches and length 18.

OpenStudy (anonymous):

ya what he said

OpenStudy (anonymous):

Thanks